# A refined Weyl character formula for comodules on $\operatorname{GL}(2,A)$

### v2

@@ 5. On the schematic normalizer in $\operatorname{GL}(2,A)$
where $\pi$ i a faithfully flat morphism and $T\cong ker(\pi)$. The affine group scheme $N/T$ satisfies the following universal property: Let $\sigma,q:N\times T\rightarrow N$ be the action map and projection map. For any morphism $f:N\rightarrow Z$ to an affine scheme $Z$ with $\sigma\circ f=q\circ f$, there is a unique morphism $\tilde{f}:N/T\rightarrow Z$ with $\tilde{f}\circ\pi=f$. Hence the affine group scheme $N/T$ is uniquely determined up to isomorphism.
Let $R:=A[x_{ij},1/D]$ where $D:=x_{11}x_{22}-x_{12}x_{21}$ is the determinant and let $G:=\operatorname{GL}(2,A):=\operatorname{Spec}(R)$. Let $I_{1}:=(x_{12},x_{21}),I_{2}:=(x_{11},x_{22})$ and $I:=I_{1}I_{2}$. It follows $I_{1}+I_{2}=(1)$ and we get an isomorphism of rings $R/I\cong R/I_{1}\oplus R/I_{2}:=R_{1}\oplus R_{2}$ with
@@
The group scheme $T$ is a normal subgroup scheme of $N$.
@@
This is a straightforward calculation. ∎
@@
Let $G:=\operatorname{GL}(2,A)$ with $A$ a commutative unital ring and let $T$ be the diagonal torus with $N\subseteq G$ the schematic normalizer. There is a faithfully flat map of group schemes $\pi:N\rightarrow N/T$ with $T\cong ker(\pi)$ giving an exact sequence of group schemes
@@
For any map $f:N\rightarrow Z$ to an affine scheme $Z$ with $f\circ\sigma=f\circ q$ there is a unique map $f^{*}:N/T\rightarrow Z$ with $f^{*}\circ\pi=f$. Hence the affine group scheme $N/T$ is uniquely determined up to isomorphism.
@@
and this is iff $f(x):=u:=a_{1}e_{1}+a_{2}e_{2}\in k[N/T]$. Hence $f\circ\sigma_{T}=f\circ q^{*}$ iff there is a map $f^{*}:C\rightarrow k[N/T]$ with $f^{*}\circ i=f$. The Theorem follows. ∎
@@
There is an isomorphism of Hopf algebras $k[N/T]\cong A[z]/(z(1-z))$, where $A[z]/(z(1-z))$ has the following structure: The comultiplication $\Delta$, coinversion $S$ and conunit $\epsilon$ is defined as follows:
@@
There is an isomorphism of groups $h_{N/T}(\mathbb{G}_{m})\cong S(2)$ where $S(2)$ is the symmetric group on two elements.
@@
Hence we recover the symmetric group on two elements as the $\mathbb{G}_{m}$-points of $h_{N/T}(-)$. Since $W_{G}\cong S(2)$, it may be we can use this to define the Weyl group in terms of the functor of points of $N/T$ in greater generality.
6. A refined Weyl character formula for $\operatorname{GL}(2,A)$.
In this section we define and calculate the Weyl group of the group schemes $\operatorname{GL}(2,A),\operatorname{SL}(2,A)$ with $A$ an arbitrary integral domain. We also introduce the refined weight space decomposition of the symmetric powers $\operatorname{Sym}^{d}(V)$ and prove in Theorem 6.4 the refined Weyl character formula. This formula implies the classical Weyl character formula.
Let $G:=\operatorname{GL}(2,A),S:=\operatorname{SL}(2,A)$ with $A$ an integral domain. Let $T_{G}\subseteq N_{G}\subseteq G$ be the diagonal torus and schematic normalizer in $G$ and let $T_{S}\subseteq N_{S}\subseteq S$ be the diagonal torus and schematic normalizer in $S$. It follows $T_{G}\cong\mathbb{G}_{m}\times\mathbb{G}_{m}$ and $T_{S}\cong\mathbb{G}_{m}$. By definition
@@
Let $G:=\operatorname{GL}(2,A)$.There is an isomorphism $W_{G}\cong S(2)$ where $S(2)$ is the symmetric group on two elements. The same result holds for $\operatorname{SL}(2,A)$.
@@
Similarly we get $s_{\beta}(\rho_{m,n})=\rho_{n,m}$. Hence the Weyl group $W_{G}\cong S(2)$ is the symmetric group on two elements. A similar calculation gives the same result for $\operatorname{SL}(2,A)$. ∎
Let $(V,\Delta)$ be the standard right comodule on $\operatorname{GL}(2,A)$. Recall the following: $V:=A\{e_{1},e_{2}\}$ and
@@
For any Schur-Weyl functor $S_{\lambda}(-)$ as introduced in [7] and for the standard right comodule $(V,\Delta)$ on $G:=\operatorname{GL}(n,\mathbf{Z})$ we may construct the comodule
@@
In a previous paper I constructed explicit examples of comodules $W_{1}\neq W_{2}$ on $\operatorname{SL}(2,\mathbf{Z})$ where the induced comodules $W_{1}\otimes\mathbf{Q}\cong W_{2}\otimes\mathbf{Q}$ are isomorphic as comodules on $\operatorname{SL}(2,\mathbf{Q})$. It is hoped the refined weight space decomposition may be helpful in the study of the problem of classifying such comodules. In general one wants for any integer $d\geq 1$ to classify finite rank torsion free comodules $W$ on $\operatorname{SL}(2,\mathbf{Z})$ such that there is an isomorphism
@@
of comodules on $\operatorname{SL}(2,\mathbf{Q})$. Here $V$ is the standard right (or left) comodule on $\operatorname{SL}(2,\mathbf{Q})$. The map $\mathbf{Z}\rightarrow\mathbf{Q}$ is flat but not faithfully, flat hence we cannot use the classical theory of faithfully flat descent for comodules to study this problem.
@@ 7. The symmetric tensors and symmetric powers
In this section we use the new methods devloped in the previous sections to give explicit examples and applications. Let $(V,\Delta)$ be the standard let comodule on $\operatorname{GL}(2,A)$ and let $\operatorname{Sym}^{2}(V),\operatorname{Sym}^{2}(V)^{*},\operatorname{S}_{2}(V),\operatorname{S}_{2}(V)^{*}$ be the symmetric powers and symmetric tensors of $V$. We study the weight space decomposition and refined weight space decomposition of the symmetric powers and symmetric tensors and prove they are isomorphic as comodules on $k[N]$ where $T\subseteq\operatorname{GL}(2,A)$ is the diagonal torus and $T\subseteq N$ is the schematic normalizer of $T$.
Let $k[G]:=A[x_{ij},1/D]$ with $G:=\operatorname{GL}(2,A):=\operatorname{Spec}(k[G]$. Let $V:=A\{e_{1},e_{2}\}$ be the free rank $2$ module on the basis $e_{1},e_{2}$ and defined the following comodule map $\Delta$:
@@
We say the pair $(V,\Delta)$ is the left standard comodule on $\operatorname{GL}(2,A)$.
@@
Let $G:=\operatorname{GL}(2,A)$ with $A$ any commutative ring, and let $N\subseteq G$ be the schematic normalizer of the diagonal torus. There is an isomorphism of $k[N]$-comodules (and $k[T]$-comodules) $\operatorname{Sym}^{2}(V)\cong\operatorname{S}_{2}(V)$. Let $G:=\operatorname{GL}(2,\mathbf{Z})$ with $\mathbf{Z}$ the ring of integers. It follows the $k[G]$-comodules $\operatorname{Sym}^{2}(V)$ and $\operatorname{S}_{2}(V)$ are not isomorphic.

### v4

@@ 5. On the schematic normalizer in $\operatorname{GL}_{2,A}$
where $\pi$ i a faithfully flat morphism and $T\cong ker(\pi)$. The affine group scheme $N/T$ satisfies the following universal property: Let $\sigma,q:N\times T\rightarrow N$ be the action map and projection map. For any morphism $f:N\rightarrow Z$ to an affine scheme $Z$ with $\sigma\circ f=q\circ f$, there is a unique morphism $\tilde{f}:N/T\rightarrow Z$ with $\tilde{f}\circ\pi=f$. Hence the affine group scheme $N/T$ is uniquely determined up to isomorphism. There is an isomorphism of group schemes
$N\cong T\times^{\sigma}N/T,$
where the right hand ”product” is the semi direct product with respect to a section $\sigma$ of $\pi$.
Let $R:=A[x_{ij},1/D]$ where $D:=x_{11}x_{22}-x_{12}x_{21}$ is the determinant and let $G:=\operatorname{GL}_{2,A}:=\operatorname{Spec}(R)$. Let $I_{1}:=(x_{12},x_{21}),I_{2}:=(x_{11},x_{22})$ and $I:=I_{1}I_{2}$. It follows $I_{1}+I_{2}=(1)$ and we get an isomorphism of rings $R/I\cong R/I_{1}\oplus R/I_{2}:=R_{1}\oplus R_{2}$ with
@@
The group scheme $T$ is a normal subgroup scheme of $N$. There is an isomorphism of group schemes $N\cong N_{G}(T)$ where $N_{G}(T)$ is the functor normliazer as defined in [8].
@@
The first claim is a straightforward calculation. The second claim is similar to Proposition 4.3. ∎
@@
The map $\rho:k[N/T]\rightarrow k[N]$ is defined as follows:
$\rho(z):=\frac{x_{11}x_{22}}{D}$
where $D:=x_{11}x_{22}-x_{12}x_{21}$ is the determinant. We get for any commutative $A$-algebra $B$ an induced map
$\pi_{B}:h_{N}(B)\rightarrow h_{N/T}(B)$
defined by $\pi_{B}(g):=\frac{a_{11}a_{22}}{D}$ where $g\in h_{N}(B)$ is the following matrix:
$g:=\begin{pmatrix}a_{11}&a_{12}\\ a_{21}&a_{22}\end{pmatrix}.$
There is a functorial section $\sigma_{B}:h_{N/T}(B)\rightarrow h_{N}(B)$ defined as follows: Let $b\in h_{N/T}(B)$ be an element with $b^{2}=b$ and define
$\sigma_{B}(b):=\begin{pmatrix}b&1-b\\ 1-b&b\end{pmatrix}.$
It follows $\sigma_{B}(b_{1}\times b_{1})=\sigma_{B}(b_{1})\sigma_{B}(b_{2})$ and $\sigma_{B}(1)=Id$. Moreover $ker(\pi_{B})=h_{T}(B)$. Hence the exact sequence
$(e)\rightarrow h_{T}(B)\rightarrow h_{N}(B)\rightarrow^{\pi_{B}}h_{N/T}(B)\rightarrow(e)$
is split by $\sigma_{B}$. It follows there is an isomorphism
$h_{N}(B)\cong h_{T}(B)\times^{\sigma_{B}}h_{N/T}(B)$
where the right hand side is the semi direct product with respect to $\sigma_{B}$. The map $\sigma_{B}$ is functorial and we get a natural transformation $\sigma:h_{N}(-)\rightarrow h_{N/T}(-)$ of functors.
@@
Let $G:=\operatorname{GL}_{2,A}$ with $A$ a commutative unital ring and let $T$ be the diagonal torus with $N\subseteq G$ the schematic normalizer. There is a faithfully flat map of group schemes $\pi:N\rightarrow N/T$ with $T\cong ker(\pi)$ giving a split exact sequence of group schemes
@@
For any map $f:N\rightarrow Z$ to an affine scheme $Z$ with $f\circ\sigma=f\circ q$ there is a unique map $f^{*}:N/T\rightarrow Z$ with $f^{*}\circ\pi=f$. Hence the affine group scheme $N/T$ is uniquely determined up to isomorphism. The section $\sigma:h_{N/T}(-)\rightarrow h_{N}(-)$ defined above gives a splitting of the sequence, hence there is an isomorphism of group schemes $N\cong T\times^{\sigma}N/T$ where the product is the semi direct product with respect to $\sigma$.
@@
and this is iff $f(x):=u:=a_{1}e_{1}+a_{2}e_{2}\in k[N/T]$. Hence $f\circ\sigma_{T}=f\circ q^{*}$ iff there is a map $f^{*}:C\rightarrow k[N/T]$ with $f^{*}\circ i=f$. One checks the section $\sigma_{B}$ defined above for any $B$ gives rise to a natural transformation of functors $\sigma:h_{N/T}(-)\rightarrow h_{N}(-)$ which is a section of the canonical map $\pi:h_{N}(-)\rightarrow h_{N/T}(-)$. The Theorem follows. ∎
@@
Let $k[S(2)]$ be the Hopf algebra of the constant group scheme (over $A$) of the symmetric group $S(2)$ on two elements. There is an isomorphism of Hopf algebras $k[S(2)]\cong A[z]/(z(1-z))$, where $A[z]/(z(1-z))$ has the following structure: The comultiplication $\Delta$, coinversion $S$ and conunit $\epsilon$ is defined as follows:
@@
Hence the group scheme $N/T$ is the constant group scheme of the Weyl group $W_{G}:=S(2)$ of $\operatorname{GL}_{2,A}$.
@@
Hence the functor normalizer $N_{G}(T)\cong T\times^{\sigma}\operatorname{Spec}(k[W_{G}])$ is the semi direct product of the diagonal torus $T\subseteq\operatorname{GL}_{2,A}$ and the constant group scheme $\operatorname{Spec}(k[W_{G}])$ of the Weyl group $W_{G}$.
6. A refined Weyl character formula for $\operatorname{GL}_{2,A}$.
In this section we define and calculate the Weyl group of the group schemes $\operatorname{GL}_{2,A},\operatorname{SL}_{2,A}$ with $A$ an arbitrary integral domain. We also introduce the refined weight space decomposition of the symmetric powers $\operatorname{Sym}^{d}(V)$ and prove in Theorem 6.4 the refined Weyl character formula. This formula implies the classical Weyl character formula.
Let $G:=\operatorname{GL}_{2,A},S:=\operatorname{SL}_{2,A}$ with $A$ an integral domain. Let $T_{G}\subseteq N_{G}\subseteq G$ be the diagonal torus and schematic normalizer in $G$ and let $T_{S}\subseteq N_{S}\subseteq S$ be the diagonal torus and schematic normalizer in $S$. It follows $T_{G}\cong\mathbb{G}_{m}\times\mathbb{G}_{m}$ and $T_{S}\cong\mathbb{G}_{m}$. By definition
@@
Let $G:=\operatorname{GL}_{2,A}$.There is an isomorphism $W_{G}\cong S(2)$ where $S(2)$ is the symmetric group on two elements. The same result holds for $\operatorname{SL}_{2,A}$.
@@
Similarly we get $s_{\beta}(\rho_{m,n})=\rho_{n,m}$. Hence the Weyl group $W_{G}\cong S(2)$ is the symmetric group on two elements. A similar calculation gives the same result for $\operatorname{SL}_{2,A}$. ∎
Let $(V,\Delta)$ be the standard right comodule on $\operatorname{GL}_{2,A}$. Recall the following: $V:=A\{e_{1},e_{2}\}$ and
@@
A relation with irreducible modules over a field of characteristic $p>0$.
Let $k$ be a field of characteristic $2$ and let $G:=\operatorname{SL}_{2,k}$. Let $V\cong k\{e_{1},e_{2}\}$ be the standard right comodule on $G$ and let $\operatorname{Sym}^{2}(V)$ be the second symmetric power of $V$. It follows the refined weight space decomposition of $\operatorname{Sym}^{2}(V)$ is as follows: Define $\operatorname{Sym}^{2}(V)^{1}:=k\{e_{1}^{2},e_{2}^{2}\}$ and $\operatorname{Sym}^{2}(V)^{2}:=k\{e_{1}e_{2}\}$. There is a direct sum decomposition of $k[N]$-comodules
$\operatorname{Sym}^{2}(V)\cong\operatorname{Sym}^{2}(V)^{1}\oplus\operatorname{Sym}^{2}(V)^{2},$
and the irreducible module $V(2)$ with highest weight $2$ is the strict submodule $V(2):=k\{e_{1}^{2},e_{2}^{2}\}\subsetneq\operatorname{Sym}^{2}(V)$. Hence there is an equality
$V(2)=\operatorname{Sym}^{2}(V)^{1}$
where $\operatorname{Sym}^{2}(V)^{1}$ is a term in the refined weight space decomposition. Hence the refined weight space decomposition contains information on irreducible modules in positive characteristic.
Example 6.7.
Irreducibility over the complex numbers.
Let $k$ be the complex number field and let $G:=\operatorname{GL}_{2,k}$. Let $(V,\Delta)$ be the standard right comodule and let $\operatorname{Sym}^{d}(V)$ be the $d$’th symmetric power with $\operatorname{Sym}^{d}(V)^{j}$ the $j$’th component in the refined weight space decomposition. The comodule
$\operatorname{Sym}^{d}(V)^{j}\rightarrow\operatorname{Sym}^{d}(V)^{j}\otimes k[N]$
is an irreducible comodule for all $j$ and $d\geq 1$.
Example 6.8.
@@
For any Schur-Weyl functor $S_{\lambda}(-)$ as introduced in [7] and for the standard right comodule $(V,\Delta)$ on $G:=\operatorname{GL}_{n,\mathbf{Z}}$ we may construct the comodule
@@
In a previous paper I constructed explicit examples of comodules $W_{1}\neq W_{2}$ on $\operatorname{SL}_{2,\mathbf{Z}}$ where the induced comodules $W_{1}\otimes\mathbf{Q}\cong W_{2}\otimes\mathbf{Q}$ are isomorphic as comodules on $\operatorname{SL}_{2,\mathbf{Q}}$. It is hoped the refined weight space decomposition may be helpful in the study of the problem of classifying such comodules. In general one wants for any integer $d\geq 1$ to classify finite rank torsion free comodules $W$ on $\operatorname{SL}_{2,A}(2,\mathbf{Z})$ such that there is an isomorphism
@@
of comodules on $\operatorname{SL}_{2,\mathbf{Q}}$. Here $V$ is the standard right (or left) comodule on $\operatorname{SL}_{2,\mathbf{Q}}$. The map $\mathbf{Z}\rightarrow\mathbf{Q}$ is flat but not faithfully, flat hence we cannot use the classical theory of faithfully flat descent for comodules to study this problem.
7. On irreducibility and the refined weight space decomposition
In this section we prove that the direct summands $\operatorname{Sym}^{d}(V)^{i}$ of the symmetric power $\operatorname{Sym}^{d}(V)$ of the standard comodule $V$ are irreducible comodules on the schematic normalizer $N$ over any field $k$.
Let $k$ be any field and let $G=\operatorname{Spec}(k[G])$ with $k[G]:=k[x_{ij},1/D]$ with $D$ the determinant. It follows for any commutative $k$-algebra $h_{G}(A)$ is the set of 2 times 2 matrices with coefficients in $A$ and determinant a unit in $A$. Let $V:=k\{e_{1},e_{2}\}$ be the standard right comodule with the following comodule map:
$\Delta_{V}:V\rightarrow k[G]\otimes_{k}V$
defined by
$\Delta_{V}(e_{1}):=x_{11}\otimes e_{1}+x_{12}\otimes e_{2},\Delta_{V}(e_{2}):=x_{21}\otimes e_{1}+x_{22}\otimes e_{2}.$
Let $k[N]:=R/I$ with variables $t_{1}:=x_{11},t_{2}:=x_{22},u_{1}:=x_{12},u_{2}:=x_{21}$ as in the previous section.
We get an induced comodule structure on the $d$’th symmetric power
$\Delta:\operatorname{Sym}^{d}(V)\rightarrow k[N]\otimes\operatorname{Sym}^{d}(V)$
defined by
$\Delta(e_{1}^{i}e_{2}^{j}):=t_{1}^{i}t_{2}^{j}\otimes e_{1}^{i}e_{2}^{j}+u_{1}^{i}u_{2}^{j}\otimes e_{1}^{j}e_{2}^{i}.$
Let $v_{i}:=e_{1}^{d-i}e_{2}^{i}$ for $i:=0,..,d$. It follows
$\operatorname{Sym}^{d}(V)^{i}\cong k\{v_{i},v_{d-i}\}.$
By construction there is for each $i=0,..,d$ an induced comodule map
$\Delta^{i}.\operatorname{Sym}^{d}(V)^{i}\rightarrow k[N]\otimes\operatorname{Sym}^{d}(V)^{i}$
defined as follows:’
$\Delta^{i}(v_{i}):=t_{1}^{d-i}t_{2}^{i}\otimes v_{i}+u_{1}^{d-i}u_{2}^{i}\otimes v_{d-i}$
and
$\Delta^{i}(v_{d-i}):=t_{1}^{i}t_{2}^{d-i}\otimes v_{d-i}+u_{1}^{i}u_{2}^{d-i}\otimes v_{i}.$
Theorem 7.1.
For each $i=0,..,d$ it follows the comodule $(\operatorname{Sym}^{d}(V)^{i},\Delta^{i})$ is an irreducible $k[N]$-comodule.
Proof.
Assume $v:=a_{1}v_{i}+a_{2}v_{d-i}$ with $a_{i}\in k$ and let $L:=kv$. It follows
$\Delta^{i}(v)=$
$(a_{1}t_{1}^{d-i}t_{2}^{i},a_{2}u_{1}^{i}u_{2}^{d-i})\otimes v_{i}+(a_{2}t_{1}^{i}t_{2}^{d-i},a_{1}u_{1}^{d-i}u_{2}^{i})\otimes v_{d-i}.$
It follows $L$ is a submodule iff there is an element $f:=(f_{1},f_{2})\in k[N]$ with
$\Delta^{i}(v)=f\otimes v=a_{1}f\otimes v_{i}+a_{2}f\otimes v_{d-i}.$
This implies the following equations:
$a_{1}t_{1}^{d-i}t_{2}^{i}=a_{1}f_{1},a_{2}u_{1}^{i}u_{2}^{d-i}=a_{1}f_{2}$
and
$a_{2}t_{1}^{i}t_{2}^{d-i}=a_{2}f_{1},a_{1}u_{1}^{d-i}u_{2}^{i}=a_{2}f_{2}.$
If $a_{1}\neq 0$ it follows $f_{1}=t_{1}^{d-i}t_{2}^{i}$ and this implies there is an equality
$a_{2}t_{1}^{i}t_{2}^{d-i}=a_{2}f_{1}=a_{2}t_{1}^{i}t_{2}^{d-i},$
hence $a_{2}=0$. This implies $a_{1}u_{1}^{d-i}u_{2}^{i}=a_{2}f_{2}=0$ hence $a_{1}=0$ a contradiction, hence $a_{1}=0$. This implies $a_{2}u_{1}^{i}u_{2}^{d-i}=a_{1}f_{2}=0$ hence $a_{2}=0$. The Theorem follows. ∎
Example 7.2.
Irreducible finite dimensional modules over a field of characteristic $p>0$.
If $G:=\operatorname{GL}(V)$ or $\operatorname{SL}(V)$ for $V:=k^{n}$ an $n$-dimensional vector space over $k$, it follows for any finite dimensional irreducible $G$-module $V(\lambda)$, there is a Weyl module $S_{\lambda}(V)$ and an inclusion $V(\lambda)\subseteq S_{\lambda}(V)$ which is strict in general. It is an open problem to calculate the character $Char(V(\lambda))$. If the terms in the refined weight space decomposition are irreducible modules over the normalizer $N$, this decomposition may be interesting in the study of such character formulas. There are canonical diract summand decompositions
$V(\lambda)\cong\oplus_{i}V(\lambda)^{i},S_{\lambda}(V)\cong\oplus_{j}S_{\lambda}(V)^{j}$
and if $S_{\lambda}(V)^{j}$ is an irreducible module, it follows the inclusion map
$f_{j}:V(\lambda)^{j}\rightarrow S_{\lambda}(V)^{j}$
is either the zero map or an isomorphism.
Note: If the schematic normalizer $N\cong T\times^{\sigma}\operatorname{Spec}(k[W_{G}])$ is the wreath product of the diagonal torus $T$ with the constand group scheme of the symmetric group $W_{G}\cong S(n)$ on n letters, we may tro to relate comodules on $N$ to comodules on $T$ and $S(n)$. We may try to relate the direct summands in the refined weight space decomposition to comodules on $T$ and $S(n)$.
@@ 8. The symmetric tensors and symmetric powers
In this section we use the new methods devloped in the previous sections to give explicit examples and applications. Let $(V,\Delta)$ be the standard let comodule on $\operatorname{GL}_{2,A}$ and let $\operatorname{Sym}^{2}(V),\operatorname{Sym}^{2}(V)^{*},\operatorname{S}_{2}(V),\operatorname{S}_{2}(V)^{*}$ be the symmetric powers and symmetric tensors of $V$. We study the weight space decomposition and refined weight space decomposition of the symmetric powers and symmetric tensors and prove they are isomorphic as comodules on $k[N]$ where $T\subseteq\operatorname{GL}_{2,A}$ is the diagonal torus and $T\subseteq N$ is the schematic normalizer of $T$.
Let $k[G]:=A[x_{ij},1/D]$ with $G:=\operatorname{GL}_{2,A}:=\operatorname{Spec}(k[G])$. Let $V:=A\{e_{1},e_{2}\}$ be the free rank $2$ module on the basis $e_{1},e_{2}$ and defined the following comodule map $\Delta$:
@@
We say the pair $(V,\Delta)$ is the left standard comodule on $\operatorname{GL}_{2,A}$.
@@
Let $G:=\operatorname{GL}_{2,A}$ with $A$ any commutative ring, and let $N\subseteq G$ be the schematic normalizer of the diagonal torus. There is an isomorphism of $k[N]$-comodules (and $k[T]$-comodules) $\operatorname{Sym}^{2}(V)\cong\operatorname{S}_{2}(V)$. Let $G:=\operatorname{GL}_{2,\mathbf{Z}}$ with $\mathbf{Z}$ the ring of integers. It follows the $k[G]$-comodules $\operatorname{Sym}^{2}(V)$ and $\operatorname{S}_{2}(V)$ are not isomorphic.